3) Steam enters a turbine from a height of 6 m, at 50m/s speed at 2MPa pressure and 350C, and after 8.5kW heat loss to the environment, it exits the turbine at 0.1Mpa pressure, with a dryness degree of 0.92, from a height of A meter at a speed of 200m/s. (A indicates a single number height)

a) Develop the expression of the 1st Law of thermodynamics by analyzing the energy of this turbine.

b) Calculate the power of the turbine if the mass of the steam entering the turbine is 1.5kg/s.

4) At E C, 30mL of coffee is forgotten in a thermos glass. If fresh coffee is added to the cooled coffee at F C, the temperature of the coffee rises to GC under room conditions. Calculate the amount of coffee added to the thermos. A two-digit number ending in 0 or 5, where E < G < F)


1
Expert's answer
2021-12-20T17:54:08-0500

"c_1V_1(G-E)=c_2V_2(F-G),"

"c_1=c_2,"

"V_2=V_1\cdot \frac{G-E}{F-G}."